OpenStudy (kise_ryouta):

Solve for x. will fan and medal x/4≥−8x a.x≤−32 b.x≤−2 c.x≥−2 d.x≥−32

OpenStudy (kise_ryouta):

@Benji

OpenStudy (welshfella):

are you sure you have the question right? because none of those options fit the inequality.

OpenStudy (welshfella):

im assuming that in 8x4 4 is an exponent

OpenStudy (kise_ryouta):

i changed it sorry for the confusion

OpenStudy (welshfella):

yea that looks better!

OpenStudy (kise_ryouta):

sorry gomen ><

OpenStudy (welshfella):

x/4 >-8x x/4 + 8x >= 0 x >= 0

OpenStudy (welshfella):

none of the options is correct

OpenStudy (mathmale):

Assuming that your "x/4≥−8x" is correct, then another way in which to write this inequality is \[\frac{ 1 }{ 4 }~x≥−8x\]

OpenStudy (mathmale):

Multiply both sides of this inequality by 4 to clear fractions:\[4(\frac{ 1 }{ 4 }~x≥−8x)\]

OpenStudy (mathmale):

This results in \[x \ge-32x\]

OpenStudy (mathmale):

Add 32x to both sides. Can you now solve the resulting inequality for x?